OpenStudy (anonymous):

I need help the polynomial 8x^3 - 343y^3 can be factored into the product of two polynomial, A*B where the degree of A is greater than the degree of B Find A and B

OpenStudy (anonymous):

turns out \(343=7^3\) so you have \[(2x)^3-(7y)^3\] which is the "difference of two cubes"

OpenStudy (anonymous):

factor difference of two cubes as \[a^3-b^3=(a-b)(a^2+ab+b^2)\] in your case \(a=2x\) and \(b=7y\)

OpenStudy (anonymous):

i got this (2x-7y)(2x^3+2x*7y+7y^3)

OpenStudy (phi):

look more carefully at sat's equation: if a= 2x and the equation has a^2 (which means a*a), then you want 2*x*2*x= 4x^2 or (2x)^2= 2^2*x^2= 4x^2 also, the middle erm 2*x*7*y can be simplified a bit: 2*7*x*y or 14xy and you must fix the last term b^2 with b=7y